Find the relative extrema using both first and second

derivative tests. f(x) = sin 2x, 0 < x < π


1
Expert's answer
2022-03-08T07:35:01-0500

f'(x)=2cos2x=0

"2x=\u03c0\/2+\u03c0n, n=0,1,2..."

"x=\u03c0\/4+\u03c0n\/2, n=0,1,2..."

f''(x)=-4sin2x

f''(π/4+πn/2)=-4sin(π/4+πn/2)=-2.84<0

So x=π/4+πn/2,n=0,1,2...x=π/4+πn/2,n=0,1,2 is minimum